# New Cold Subdwarf Discoveries from Backyard Worlds and a Metallicity Classification System for T Subdwarfs

### v1

@@ 3.6 Additional Spectra
J2218+1146ccAlso identified as a T subdwarf candidate by our selection criteria. | J221859.41+114642.9 | 17.960.03 | T7p | Keck/NIRES | 51 | [22] |
@@ 4.4 Spectral Model Fits
1637 | 5.47 | 0.05 | 0.21bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 4 | SAND | |||||
@@
1536 | 5.72 | 0.03 | 0.08bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 6 | SAND | |||||
@@
1663 | 5.47 | 0.01 | 0.08bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 2 | SAND | |||||
@@
1803 | 5.57 | 0.20 | 0.06bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 8 | SAND | |||||
@@
1565 | 5.35 | 0.69 | 0.20bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 9 | SAND | |||||
@@
1611 | 5.59 | 0.35 | 0.34bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 8 | SAND | |||||
@@
1520 | 5.63 | 0.74 | 0.38bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 4 | SAND | |||||
@@
2011 | 4.80 | 0.56 | 0.20bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 7 | SAND | |||||
@@
727 | 5.03 | 0.61 | 0.18bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 80 | Settl | |||||
@@
1859 | 4.80 | 0.15 | 0.10bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 14 | Settl | |||||
@@
1488 | 5.65 | 0.44 | 0.18bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.063 | 12 | SAND | ||||
@@
1718 | 5.05 | 1.53 | 0.33bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.077 | 9 | SAND | ||||
@@
2181 | 5.20 | 1.13 | 0.07bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.056 | 6 | Dusty | ||||
@@
1594 | 5.49 | 0.46 | 0.29bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.063 | 65 | SAND | ||||
@@
1852 | 5.53 | 0.17 | 0.05bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.099 | 46 | SAND | ||||
@@
1695 | 5.59 | 0.01 | 0.17bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 10 | SAND | |||||
@@
1870 | 5.50 | 0.19 | 0.09bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.062 | 30 | SAND | ||||
@@
1633 | 5.45 | 0.42 | 0.35bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 1 | SAND | |||||
@@
1639 | 5.51 | 0.50 | 0.22bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 1 | SAND | |||||
@@
869 | 5.87 | 1.27 | 0.34bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.081 | 9 | SAND | ||||
J21056235 | d/sdT2: | Flamingos | 1282 | 4.72 | 0.32 | 0.66 | 10 | 65 | LOWZ |
@@ 7.1 Effective Temperature Trends for T Subdwarfs
J15341043 | (e)sdY: | 23.40.3 | 0.56ddModel fit produces non-physically small radius (R 0.05 R⊙), suggesting an overestimated . | 22.9 | 31432 | 15.090.19 | 6.26eeUsing the BC value for = 500 K and [M/H] = 1.5. | 21.35 | 44745 | 38167 |

### v2

@@ 3.6 Additional Spectra
| J2218+1146ccAlso identified as a T subdwarf candidate by our selection criteria. | J221859.41+114642.9 | 17.960.03 | T7p | Keck/NIRES | 51 | [22] | |
@@ 4.4 Spectral Model Fits
| 1637 | 5.47 | 0.05 | 0.21bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 4 | SAND | |||||
@@
| 1536 | 5.72 | 0.03 | 0.08bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 6 | SAND | |||||
@@
| 1663 | 5.47 | 0.01 | 0.08bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 2 | SAND | |||||
@@
| 1803 | 5.57 | 0.20 | 0.06bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 8 | SAND | |||||
@@
| 1565 | 5.35 | 0.69 | 0.20bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 9 | SAND | |||||
@@
| 1611 | 5.59 | 0.35 | 0.34bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 8 | SAND | |||||
@@
| 1520 | 5.63 | 0.74 | 0.38bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 4 | SAND | |||||
@@
| 2011 | 4.80 | 0.56 | 0.20bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 7 | SAND | |||||
@@
| 727 | 5.03 | 0.61 | 0.18bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 80 | Settl | |||||
@@
| 1859 | 4.80 | 0.15 | 0.10bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 14 | Settl | |||||
@@
| 1488 | 5.65 | 0.44 | 0.18bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.063 | 12 | SAND | ||||
@@
| 1718 | 5.05 | 1.53 | 0.33bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.077 | 9 | SAND | ||||
@@
| 2181 | 5.20 | 1.13 | 0.07bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.056 | 6 | Dusty | ||||
@@
| 1594 | 5.49 | 0.46 | 0.29bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.063 | 65 | SAND | ||||
@@
| 1852 | 5.53 | 0.17 | 0.05bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.099 | 46 | SAND | ||||
@@
| 1695 | 5.59 | 0.01 | 0.17bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 10 | SAND | |||||
@@
| 1870 | 5.50 | 0.19 | 0.09bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.062 | 30 | SAND | ||||
@@
| 1633 | 5.45 | 0.42 | 0.35bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 1 | SAND | |||||
@@
| 1639 | 5.51 | 0.50 | 0.22bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 1 | SAND | |||||
@@
| 869 | 5.87 | 1.27 | 0.34bbAlpha element enrichment [/Fe] relative to the Sun for Settl and Dusty models. | 0.081 | 9 | SAND | ||||
| J21056235 | d/sdT2: | Flamingos | 1282 | 4.72 | 0.32 | 0.66 | 10 | 65 | LOWZ | |
@@ 7.1 Effective Temperature Trends for T Subdwarfs
| J15341043 | (e)sdY: | 23.40.3 | 0.56ddModel fit produces non-physically small radius (R 0.05 R⊙), suggesting an overestimated . | 22.9 | 31432 | 15.090.19 | 6.26eeUsing the BC value for = 500 K and [M/H] = 1.5. | 21.35 | 44745 | 38167 | |